OpenStudy (anonymous):

log_{2}(2-x)-2=log_{2} (x+8)

OpenStudy (anonymous):

\[\log_{2}(2-x)-2=\log_{2} (x+8)\]

OpenStudy (anonymous):

step by step please

jimthompson5910 (jim_thompson5910):

\[\large \log_{2}(2-x)-2=\log_{2} (x+8)\] \[\large \log_{2}(2-x)-\log_{2}(4)=\log_{2} (x+8)\] \[\large \log_{2}\left(\frac{2-x}{4}\right)=\log_{2} (x+8)\] \[\large \frac{2-x}{4}=x+8\] \[\large 2-x=4(x+8)\] \[\large 2-x=4x+32\] \[\large 2-32=4x+x\] \[\large -30=5x\] \[\large \frac{-30}{5}=x\] \[\large -6=x\] \[\large x=-6\]

jimthompson5910 (jim_thompson5910):

notes: I converted 2 to \[\log_2(4)\] because I wanted everything inside a log with base 2

OpenStudy (anonymous):

so 2 = \[\log_{2}4 \] ?

jimthompson5910 (jim_thompson5910):

\[\large \log_{2}(4)=\log_{2}(2^2)=2\log_{2}(2)=2(1)=2\] So \[\large \log_{2}(4)=2\]

OpenStudy (anonymous):

thank you again!